OpenStudy (anonymous):

I just need to check my answers! 1. The lengths of the sides of a rhombus and one of its diagonals are each 10m. What is the area of the rhombus? a. 100√3m² b. 50√3m² c. 25√3m² d. 12.5√3m² My answer is B. 2. The area of a regular decagon is 50 cm². What is the area of a larger regular decagon with sides four times the sides of the smaller decagon? a. 200 cm² b. 500 cm² c. 800 cm² d. 2000 cm² I have A. Thanks! x

OpenStudy (raden):

the first one is right. but for Q2, that should be the larger area = 4^2 * the smaller area = 16 * 50 = 800

OpenStudy (anonymous):

Thanks so much!

OpenStudy (raden):

you're welcome